# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ Abstract
This paper constructs a full exceptional collection in the bounded derived category of coherent sheaves on a generalized flag scheme $G/P$ over $\mathbb{Z}$, by means of a $G$-linear semiorthogonal decomposition of the bounded derived category of noetherian representations of $P$. The ensuing semiorthogonal decompositions are compatible with the Bruhat order on cosets of the Weyl group of $P$ in the Weyl group of $G$. The construction builds upon the foundational results on $B$-modules from the works of Mathieu, Polo, and van der Kallen, and upon properties of the Steinberg basis of the $T$-equivariant ${\sf K}$-theory of $G/B$.
@@
flag variety, exceptional collection, $B$-module, equivariant $K$-theory
@@
Acknowledgements
A.S. is grateful to Henning Krause for his unwavering support of this work and for his and Charles Vial’s joint invitation to the BIREP group of the University of Bielefeld for a long-term visit in 2023-2024. A.S. work was supported by the Deutsche Forschungsgemeinschaft (SFB-TRR 358/1 2023 - 491392403). He also acknowledges the hospitality of the Mathematical Institute of the University of Utrecht during his visit in May 2024.
@@ 1.1 The context
Let $\mathbb{G}$ be a split semi-simple simply connected Chevalley group scheme and $\mathbb{B}$ a split Borel subgroup scheme. Associated to it is the flag scheme $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$. This paper constructs full exceptional collections in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$. More generally, let $\mathbb{P}$ be a parabolic subgroup scheme containing $\mathbb{B}$. We also construct full exceptional collections in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{P}$ over $\mathbb{Z}$.
@@
The problem of finding full exceptional collections on flag varieties and, more generally, on quotients of $\mathbb{G}$ modulo a parabolic subgroup ${\mathbb{P}}$ has a long history; we refer the reader to the paper [22] and the references therein. In this non-technical part of the introduction we emphasise some of the features that make the approach of the present paper different from many of the existing constructions.
@@
Given a smooth proper Noetherian scheme over a field ${\sf k}$, a full exceptional collection in the bounded derived category of coherent sheaves $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ is, informally speaking, a way to break up $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ into elementary pieces, each equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{vect}\textrm{-}{\sf k}}\nolimits)$ which are glued to each other in a non-trivial way. Upon decategorification, a full exceptional collection gives rise to a basis of the Grothendieck group ${{\sf K}}^{0}(X)$. If one is interested only in the ${\sf K}$-theoretic information, for many schemes of interest it is easy to produce a basis of ${{\sf K}}^{0}(X)$: for instance, if a scheme $X$ has a ${\mathbb{G}}_{m}$-action with isolated fixed points then by Białynicki-Birula’s theorem the scheme $X$ has a stratification into locally closed subschemes such that the structure sheaves of their closures are a basis of ${{\sf K}}^{0}(X)$. Flag schemes $\mathbb{G}/\mathbb{B}$ furnish a classical example of such a stratification: in this case, the closures of algebraic cells of the Białynicki-Birula’s stratification are Schubert schemes via the Bruhat decomposition.
@@
Given a scheme as above and assuming it has a natural geometric basis of the group ${{\sf K}}^{0}(X)$, one can ask whether such a basis can be lifted to a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. Speaking less loosely, a full exceptional collection must satisfy two conditions: it should generate the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ in a suitable sense and some cohomological vanishing should hold. Generation doesn’t pose a problem: for schemes as above, i.e. having an algebraic stratification into locally closed subschemes, which are isomorphic to affine spaces, the direct sum of structure sheaves of their closures is a classical generator of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. But there is no chance that for such a set of generators the cohomological vanishing (known as “semiorthogonality” condition) would hold. On the other hand, there is more to flag schemes that one would want to take into account: there is a partial order on Schubert schemes (the Bruhat order) and the order on the sought-for full exceptional collection is expected to be compatible with the Bruhat order on Schubert schemes. Thus, what one is after is finding a correct “lift” of the natural ${\sf K}$-theoretic basis to the derived category. In the present paper, we provide such a lift. As explained below, there are two major inputs into our construction: one is representation-theoretic and comes from the highest weight category structures on the category of rational representations of the Borel group scheme $\mathbb{B}$, and the other one is combinatorial and comes from the distinguished basis of the ($\mathbb{T}$-equivariant) ${\sf K}$-group of the flag scheme $\mathbb{G}/\mathbb{B}$, called the Steinberg basis.
@@
There is a basic way to relate algebraic representations of $\mathbb{B}$ to $\mathbb{G}$-equivariant vector bundles on $\mathbb{G}/\mathbb{B}$ via the “associated sheaf” construction. Note that the latter construction was the starting point in [34] (see Section 1.3 below for more on that). The category ${\mathrm{rep}}(\mathbb{G})$ of rational modules over a reductive group scheme $\mathbb{G}$ has a highest weight category structure, [13], [17], and it is also such a structure that one expects to manifest itself through conjectural full exceptional collections. In this respect, a model example of such semiorthogonal decompositions/full exceptional collections is given by a theorem of Efimov’s [20] that solves the problem for the Grassmannians $\sf Gr_{k,n}$ in the best possible way. Precisely, one has (cf. also [8]):
Theorem 1.1 ([20, Theorem 1.8]).
@@
The above theorem 1.1 and, more generally, results of [20] and [8] rely considerably on representation theory of the full linear group $GL_{n}$ and on the highest weight category structure on ${\mathrm{rep}}(GL_{n})$. Remarkably, highest weight category structures have also been known to exist on the category ${\mathrm{rep}}(B)$ of rational $B$-modules for the Borel subgroup $B\subset G$ through the works [35], [41], [46] and [47]. It seems, however, that the category ${\mathrm{rep}}(B)$ as a highest weight category has not yet received the full attention that it undoubtedly deserves. The following quote from Stephen Donkin, [19, Remark 4.7] has served as a particular impetus for the first named author: “In fact ${\mathsf{k}}[{B}]$ is a quasi–hereditary coalgebra. This follows from van der Kallen’s paper [46]. This is a deep and sophisticated work which generalizes the above (restriction gives a full embedding of the category of $G$-modules into the category of $B$-modules) and whose consequences, to the best of my knowledge, have so far not been investigated or exploited.”
In the present paper, we demonstrate the force of highest weight category structures on the category ${\mathrm{Rep}}(\mathbb{B})$ and, in accordance with Donkin’s suggestion, develop one of its consequences constructing full exceptional collections on flag schemes $\mathbb{G}/\mathbb{B}$. In that context, the two main theorems are
@@
The category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ has a $G$-linear semiorthogonal decomposition
@@
with respect to the total order $\prec$ on the Weyl group $W$. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Let $v,w$ denote elements of $W$. Then there are objects ${\mathcal{X}}_{v}\in\mathcal{D}$ such that
@@
For more details, see Section 11 and Section 12, respectively. And see section 13 for the treatment of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$, $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/P))$.
By its very design, highest weight category structure on an abelian category gives rise to two distinguished collections of objects (standard and costandard objects) that satisfy the ${\mathrm{Ext}}$-vanishing close to the one that is required for exceptional collections. It would therefore be reasonable to start off with one of such collections in ${\mathrm{Rep}}(\mathbb{B})$, convert them into a collection of equivariant vector bundles, and try to ensure the cohomological vanishing conditions in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits({\mathbb{G}}/{\mathbb{B}}))$. It turns out that this doesn’t work on the nose, but it does work with the input provided by the combinatorics of the Steinberg basis. Coming back to the problem of lifts of $\sf K$-theoretic classes to the derived category level, it is important to mention the classical works [5] and [16] that compute the singular cohomology and, respectively, the $\sf K$-theory of flag schemes $\mathbb{G}/\mathbb{B}$ in terms of natural operators called nowadays BGG-Demazure operators. These operators, in different guises, will appear throughout the paper: one of the highest weight category structures on ${\mathrm{rep}}(B)$ is defined in terms of those, see Section 2. We now proceed to the second part of the introduction in which we outline the main steps of the argument.
@@ 1.2 Outline of the proof
Evidence for the principal construction of this paper comes from semiorthogonal decompositions for flag varieties of rank two groups over ${\mathbb{Z}}[\frac{1}{6}]$ established in [42]. We provide more detail of the rank two case in Section 14 below.
@@
For the purposes of introduction, assume for simplicity that $\mathbb{P}=\mathbb{B}$ and that we are working over a field ${\sf k}$ whose characteristic is not too small. [^1] We actually sometimes worked over ${\mathbb{C}}$, sometimes over ${\mathbb{Z}}$.; thus, $B$ and $G$ denote the base change of group schemes $\mathbb{B}$ and $\mathbb{G}$ along the morphism ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$, [18, Section 2.6]. Given a finite dimensional $B$-module $M$, let $\mathcal{L}(M)$ denote the associated (locally free) coherent sheaf on $G/B$. The examples in rank two have full exceptional collections consisting of locally free sheaves that are in fact obtained via the forgetful functor ${\mathrm{Rep}}(B)\mapsto{\mathop{\sf{Coh}}\nolimits}({G/B})$ and therefore have an underlying $B$-module.
@@
With those rank two examples in hand, it has become apparent that the $B$-module theory developed in [47] would be vital in generalizing the approach to arbitrary rank. One way to summarize the results of loc.cit. is saying that the abelian category ${\mathrm{Rep}}(B)$ has two highest weight category structures; the costandard objects with respect to one structure are denoted $P(\lambda)$, $\lambda\in X(T)$ (where $X(T)$ is the weight lattice of a maximal torus) and are related to excellent filtrations on $B$-modules, see Section 2.7. The costandard objects with respect to another highest weight category structure are denoted $Q(\lambda)$, $\lambda\in X(T)$ and are related to relative Schubert filtrations on $B$-modules, see Section 2.3. It is these modules that will play a pivotal role in the construction of full exceptional collections on flag schemes. We refer the reader to [46] for the highest weight category approach to ${\mathrm{Rep}}(B)$; in [47] those structures are somewhat hidden in the background.

### v3

# Highest weight category structures on Rep($\mathbf{B}$) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

@@ Abstract
Given a split reductive Chevalley group scheme $\mathbb{G}$ over $\mathbb{Z}$ and a parabolic subgroup scheme $\mathbb{P}\subset\mathbb{G}$, this paper constructs $\mathbb{G}$-linear semiorthogonal decompositions of the bounded derived category of noetherian representations of $\mathbb{P}$ with each semiorthogonal component being equivalent to the bounded derived category of noetherian representations of $\mathbb{G}$. The $\mathbb{G}$-linear semiorthogonal decompositions in question are compatible with the Bruhat order on cosets of the Weyl group of $\mathbb{P}$ in the Weyl group of $\mathbb{G}$. Their construction builds upon the foundational results on $\mathbb{B}$-modules from the works of Mathieu, Polo, and van der Kallen, and upon properties of the Steinberg basis of the $\mathbb{T}$-equivariant ${\sf K}$-theory of $\mathbb{G}/\mathbb{B}$. As a corollary, we obtain full exceptional collections in the bounded derived category of coherent sheaves on generalized flag schemes $\mathbb{G}/\mathbb{P}$ over $\mathbb{Z}$.
@@
$B$-modules, equivariant $K$-theory, Steinberg basis, flag variety, derived category, semiorthogonal decompositions, exceptional collections
@@ 1.1 The context
Let $\mathbb{G}$ be a split semi-simple simply connected Chevalley group scheme and $\mathbb{B}\subset\mathbb{G}$ a Borel subgroup scheme. Associated to it is the flag scheme $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$. More generally, let $\mathbb{P}\subset\mathbb{G}$ be a parabolic subgroup scheme containing $\mathbb{B}$. Generalized flag varietes for the group scheme $\mathbb{G}$ are the quotient schemes $\mathbb{G}/\mathbb{P}$.
The goal of this paper is to initiate a study of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbb{B}))$, the bounded derived category of noetherian representations of $\mathbb{B}$ (and, more generally, of the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbb{P}))$). Our main results, which are Theorems 11.6 and 13.3, construct some particular $\mathbb{G}$-linear semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbb{B}))$ (resp., $\mathbb{G}$-linear semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbb{P}))$) that can be considered as categorifications of the classical results [8] and [19]. As a corollary to those theorems, we obtain in Theorem 12.1 (resp., in Theorem 13.5) full exceptional collections in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{B}$ over $\mathbb{Z}$ (resp., in the bounded derived category of coherent sheaves on $\mathbb{G}/\mathbb{P}$).
We believe that the categorical decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbb{B}))$ will have further applications in geometric representation theory (see Section 1.3 below). Chronologically, we arrived at Theorems 11.6 and 13.3 starting off with the question whether full exceptional collections on generalized flag varietes can be obtained using representation theory of a Borel subgroup scheme $\mathbb{B}$. We now explain this path in a greater detail.
@@
The study of exceptional collections on generalized flag varieties has a long history; more recently, the paper [37] set out an approach that influenced many papers on the subject. For a very recent and comprehensive survey of the works that followed loc.cit., we refer the reader to [25] and the references therein. In this non-technical part of the introduction we emphasise some of the features that make the approach of the present paper different from the previous constructions.
@@
Given a smooth proper Noetherian scheme over a field ${\sf k}$, a full exceptional collection in the bounded derived category of coherent sheaves $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ is, informally speaking, a way to break up $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ into elementary pieces, each equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{vect}\textrm{-}{\sf k}}\nolimits)$ which are glued to each other in a non-trivial way. Upon decategorification, a full exceptional collection gives rise to a basis of the Grothendieck group ${{\sf K}}^{0}(X)$. If one is interested only in the ${\sf K}$-theoretic information, for many schemes of interest it is easy to produce a basis of ${{\sf K}}^{0}(X)$: for instance, if a scheme $X$ has a ${\mathbb{G}}_{m}$-action with isolated fixed points then by Białynicki-Birula’s theorem $X$ has a stratification into locally closed subschemes such that the structure sheaves of their closures are a basis of ${{\sf K}}^{0}(X)$. Flag schemes $\mathbb{G}/\mathbb{B}$ furnish a classical example of such a stratification: in this case, the closures of algebraic cells of the Białynicki-Birula’s stratification are Schubert schemes via the Bruhat decomposition.
@@
Given a scheme $X$ as above and assuming it has a natural geometric basis of the group ${{\sf K}}^{0}(X)$, one can ask whether such a basis can be lifted to a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. Speaking less loosely, a full exceptional collection must satisfy two conditions: it should generate the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$ in a suitable sense and some cohomological vanishing should hold. Generation doesn’t pose a problem: for schemes as above, i.e. having an algebraic stratification into locally closed subschemes, which are isomorphic to affine spaces, the direct sum of structure sheaves of their closures is a classical generator of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(X))$. But there is no chance that for such a set of generators the cohomological vanishing (known as “semiorthogonality” condition) would hold. On the other hand, there is more to flag schemes that one would want to take into account: there is a partial order on Schubert schemes (the Bruhat order) and the order on the sought-for full exceptional collection is expected to be compatible with the Bruhat order on Schubert schemes. Thus, what one is after is finding a correct “lift” of the natural ${\sf K}$-theoretic basis to the derived category. In the present paper, we provide such a lift. As explained below, there are two major inputs into our approach: one is representation-theoretic and comes from the highest weight category structures on the category of rational representations of the Borel group scheme $\mathbb{B}$, and the other one is combinatorial and comes from the distinguished basis of the ($\mathbb{\mathbf{T}}$-equivariant) ${\sf K}$-group of the flag scheme $\mathbb{G}/\mathbb{B}$, called the Steinberg basis, [46].
@@
There is a basic way to relate algebraic representations of $\mathbb{B}$ to $\mathbb{G}$-equivariant vector bundles on $\mathbb{G}/\mathbb{B}$ via the “associated sheaf” construction. Note that the latter construction was the starting point in [37] (see Section 1.3 below for more on that). Let $\mathbf{B}$ (resp., $\mathbf{G}$) denote the group schemes over ${\sf k}$ obtained by base change from $\mathbb{B}$ (resp., $\mathbb{G}$) along ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$, where ${\sf k}$ is a field. The category ${\mathrm{rep}}(\mathbf{G})$ of rational modules over the group scheme $\mathbf{G}$ that are finite-dimensional over ${\sf k}$ has a highest weight category structure, [16], [20], and it is these structures that are expected to manifest themselves through conjectural full exceptional collections. In this respect, a model example of such semiorthogonal decompositions/full exceptional collections is given by a theorem of Efimov’s [23] that solves the problem for the Grassmannians $\sf Gr_{k,n}$ in the best possible way. Precisely, one has (cf. also [11] for a different approach over a field ${\sf k}$):
Theorem 1.1 ([23, Theorem 1.8]).
@@
The above theorem 1.1 and, more generally, results of [23] and [11] rely considerably on representation theory of the full linear group $GL_{n}$ and on the highest weight category structure on ${\mathrm{rep}}(GL_{n})$. Remarkably, through the works [38], [43], [48] and [49], highest weight category structures have also been known to exist on the category ${\mathrm{rep}}(\mathbf{B})$ for a Borel subgroup $\mathbf{B}\subset\mathbf{G}$. Here ${\mathrm{rep}}(\mathbf{B})$ stands for the category of rational $\mathbf{B}$-modules that are finite-dimensional over ${\sf k}$. It seems, however, that the category ${\mathrm{rep}}(\mathbf{B})$ as a highest weight category has not yet received the full attention that it undoubtedly deserves. The following quote from Stephen Donkin, [22, Remark 4.7] had served as a particular impetus for the first named author: “In fact ${\mathsf{k}}[{\mathbf{B}}]$ is a quasi–hereditary coalgebra. This follows from van der Kallen’s paper [48]. This is a deep and sophisticated work which generalizes the above (restriction gives a full embedding of the category of $\mathbf{G}$-modules into the category of $\mathbf{B}$-modules) and whose consequences, to the best of my knowledge, have so far not been investigated or exploited.”
In the present paper, we demonstrate the force of highest weight category structures on the category ${\mathrm{Rep}}(\mathbf{B})$ and, in accordance with Donkin’s suggestion, show some of its consequences. Our two main theorems are:
@@
The category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ has a $\mathbf{G}$-linear semiorthogonal decomposition
@@
with respect to a total order $\prec$ on the Weyl group $W$ that refines the Bruhat order. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
As it will become clear, the case of Borel subgroups is the most important. With all the setup developed for proving Theorem 1.2, we are in a position to prove the parabolic version of it. Let $\mathbf{P}$ be a parabolic subgroup containing $\mathbf{B}$. (If ${\sf k}={\mathbb{Z}}$, then $\mathbf{P}$ means $\mathbb{P}$.) Let $W_{\mathbf{P}}$ be the parabolic Weyl group corresponding to $\mathbf{P}$, and $W^{\mathbf{P}}$ be the set of minimal coset representatives of $W/W_{\mathbf{P}}$. Let $\prec_{\mathbf{P}}$ denote the restriction to $W^{\mathbf{P}}$ of the chosen total order $\prec$ on $W$ from Theorem 1.2 above. Then:
@@
The category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{P}}))$ has a $\mathbf{G}$-linear semiorthogonal decomposition
$\mathcal{D}=\langle{\hat{\sf X}}_{v}\rangle_{v\in W^{\mathbf{P}}}$ (2)
with respect to the order $\prec_{\mathbf{P}}$ on $W^{\mathbf{P}}$. Each subcategory ${\hat{\sf X}}_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
Theorems 11.6 and 12.1 have their counterparts over $\mathbb{Z}$ if one replaces $\mathbf{G}$ (resp., $\mathbf{B}$) by $\mathbb{G}$ (resp., by $\mathbb{B}$). For more details, see Section 11 and Section 12, respectively.
Theorems 1.2 and 1.3 have some immediate applications.
Theorem 1.4 (Theorem 12.1).
Let $\prec$ be the same total order on $W$ as in Theorem 1.2, and let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$. Let $v,w$ denote elements of $W$. Then there are objects ${\mathcal{X}}_{v}\in\mathcal{D}$ such that
@@
The next Theorem 1.5 is in the same position with respect to the previous Theorem 1.4 as Theorem 1.3 is with respect to Theorem 1.2:
Theorem 1.5 (Theorem 13.5).
Let $\prec_{\mathbf{P}}$ be the same total order on $W^{\mathbf{P}}$ as in Theorem 1.3, and let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P}))$. Let $v,w\in W^{\mathbf{P}}$. Then there are objects $\hat{\mathcal{X}}_{v}\in\mathcal{D}$ such that
  1. 1.
  2. 2.
    If $w\succ_{\mathbf{P}}v$ then $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(\hat{\mathcal{X}}_{v},\hat{\mathcal{X}}_{w}[i])=0$ for all $i$.
  3. 3.
    The triangulated hull of $\{\hat{\mathcal{X}}_{v}\mid v\in W^{\mathbf{P}}\}$ is $\mathcal{D}$.
In other words, the collection of objects $(\hat{\mathcal{X}}_{v})_{v\in W^{\mathbf{P}}}$ is a full exceptional collection in $\mathcal{D}$.
In Sections 12 and 13 we explain how Theorem 1.4 (resp., Theorem 1.5) follows immediately from Theorem 1.2 (resp., Theorem 1.3). Still, our line of thought worked backwards and we conclude this section with a few preliminary remarks explaining our logic. By its very design, the highest weight category structure on an abelian category gives rise to two distinguished collections of objects (standard and costandard objects) that satisfy the ${\mathrm{Ext}}$-vanishing close to the one that is required for exceptional collections. It would therefore be reasonable to start off with one of such collections in ${\mathrm{Rep}}(\mathbf{B})$, convert them into a collection of equivariant vector bundles, and try to ensure the cohomological vanishing conditions in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits({\mathbf{G}}/{\mathbf{B}}))$. It turns out that this doesn’t work on the nose, but it does work with the input provided by the combinatorics of the Steinberg basis, [46]. Coming back to the problem of lifts of $\sf K$-theoretic classes to the derived category level, we have already mentioned the classical works [8] and [19] that compute the singular cohomology and, respectively, the $\sf K$-theory of flag varieties $\mathbf{G}/\mathbf{B}$ in terms of natural operators called nowadays BGG-Demazure operators. These operators, in different guises, will appear throughout the paper: one of the highest weight category structures on ${\mathrm{rep}}(\mathbf{B})$ is defined in terms of those, see Section 2. We now proceed to the second part of the introduction in which we outline the main steps of the argument.
@@ 1.2 Outline of the proof
Evidence for the principal construction of this paper comes from special full exceptional collections on flag varieties of rank two groups over ${\mathbb{Z}}[\frac{1}{6}]$, the proofs of which were sketched in [44]. We provide the complete details of the rank two case in Section 14. For the purposes of introduction, assume for simplicity that $\mathbb{P}=\mathbb{B}$ and that we are working over a field ${\sf k}$ whose characteristic is not too small. Thus, $\mathbf{B}$ and $\mathbf{G}$ denote the base change of the group schemes $\mathbb{B}$ and $\mathbb{G}$ along the morphism ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$, [21, Section 2.6]. Given a finite dimensional $\mathbf{B}$-module $M$, let $\mathcal{L}(M)$ denote the associated (locally free) coherent sheaf on $\mathbf{G}/\mathbf{B}$. The full exceptional collections from Section 14 are given by locally free sheaves that are obtained via the forgetful functor ${\mathrm{Rep}}(\mathbf{B})\rightarrow{\mathop{\sf{Coh}}\nolimits}({\mathbf{G}/\mathbf{B}})$ and therefore have underlying $\mathbf{B}$-module structure. That the terms of those collections are $\mathbf{G}$-equivariant vector bundles is in accordance with a general fact on exceptional objects on smooth projective varieties acted upon by a linear algebraic group, [42, Lemma 2.2]. More importantly, the collections in question are obtained with the help of Demazure functors (see Section 14.4). That points out to the fact that the highest weight category structures on ${\mathrm{rep}}(\mathbf{B})$ can be relevant: as was already mentioned, one of those structures is defined using Demazure operators, see Section 2.5.
@@
With those rank two examples in hand, it has become apparent that the $\mathbf{B}$-module theory developed in [48] would be vital in generalizing the approach to arbitrary rank. One way to summarize the results of loc.cit. is saying that the abelian category ${\mathrm{rep}}(\mathbf{B})$ has two highest weight category structures; recall that simple modules in ${\mathrm{rep}}(\mathbf{B})$ are parametrized by $X(\mathbf{T})$, the weight lattice of a maximal torus. Now the costandard objects with respect to one structure, which is defined by the excellent order on $X(\mathbf{T})$, (Definition 2.4) are denoted $P(\lambda)$, $\lambda\in X(\mathbf{T})$ and are related to excellent filtrations on $\mathbf{B}$-modules, see Section 2.5. The costandard objects with respect to another highest weight category structure, which is defined by the antipodal excellent order on $X(\mathbf{T})$, (Definition 2.5) are denoted $Q(\lambda)$, $\lambda\in X(\mathbf{T})$ and are related to relative Schubert filtrations on $\mathbf{B}$-modules, see Section 2.6. It is these modules that will play a pivotal role in the construction of full exceptional collections on generalized flag schemes. We refer the reader to [48] for the highest weight category approach to ${\mathrm{rep}}(\mathbf{B})$; in [49] those structures are somewhat hidden in the background. The lectures [49] are devoted to the breakthrough paper [39] by Mathieu; that breakthrough will be used in the guise of Corollary 2.21.